solve: (0.3)^(x+1)=0.7^x
Answer by nerdybill(7384)

(0.3)^(x+1)=0.7^x
take the log of both sides:
log[(0.3)^(x+1)] = log[0.7^x]
(x+1)*log[0.3] = x*log[0.7]
x*log[0.3]+log[0.3] = x*log[0.7]
log[0.3] = x*log[0.7]- x*log[0.3]
log[0.3] = x(log[0.7]- log[0.3])
log[0.3]/(log[0.7]- log[0.3]) = x
-1.421 = x